Sergio recently graduated with a degree in financial counseling and planning and now works for a local non-profit agency as a family financial counselor earning $45,600 per year with about $35,300 in take-home pay. He rents an apartment for $614 per month. While in school, he accumulated about $26,946 in student loan debt on which he pays $228 per month. During his last semester in school, he had an internship in a 12-county region for extension and outreach. He used his credit card for travel expenses and has a current debt on the account of $5,700. He has been making the minimum payment on the account of about $200 a month. Calculate Sergio’s debt-to-income ratio. (5 points) Remember to convert all dollar figures used to annual or monthly.

Debt payments

Gross income

Debt-to-income ratio
Answer #1

Answer:

Gross monthly income = 45600 /12 = $3,800

Monthly debt payments = 614 + 228 + 200 = $1042

Debt-to-income ratio = 1042 /3800 = 0.2742 or 27.42%

Debt-to-income ratio = 0.2742 or 27.42%